OpenStudy (anonymous):

integral of |3^x-2^x|dx?

OpenStudy (anonymous):

as x goes from -1 to 1? sorry i forgot to put that in the question.

OpenStudy (shubhamsrg):

lets compare 3^x and 2^x for 3^x > 2^x x ln 3 > x ln 2 x(ln 3 - ln 2)>0 thus for x>0, 3^x > 2^x or 3^x - 2^x is positive for x=0 , 3^x -2^x =0 for x<0, 3^x - 2^x is -ve so our inetgral can be broken like this : integral(|3^x-2^x|) [-1,1] = integral(2^x - 3^x) [-1,0) + 0 + integral ( 3^x -2^x) (0,1] can you simplify further?

OpenStudy (anonymous):

\[f(x) = \left(\begin{matrix}-(3^x - 2^x), for (3^x - 2^x) < 0 \\ (3^x - 2^x), for(3^x-2^x) \ge 0\end{matrix}\right)\]